OpenStudy (anonymous):

Solve for x: -3|x - 3| = -6 x = 1, x = -1 x = 1, x = 5 x = 0, x = 5 No solutions

OpenStudy (anonymous):

any one?

OpenStudy (johnweldon1993):

First things first, you need to get the absolute value by itself, So lets divide both sides by -3 \[\large |x-3| = 2\] Now this breaks into 2 equations \(\large x - 3 = 2\) and \(\large x - 3 = -2\) Solve those 2 equations for 'x'

OpenStudy (anonymous):

im still confused

OpenStudy (johnweldon1993):

On what? :)

OpenStudy (johnweldon1993):

So we start with \[\large -3|x - 3| = -6\] The first goal here, is to get the absolute value by itself...and we see it is being multiplied by a -3...so we need to divide both side of the equation by -3 to cancel that out \[\large \frac{-3|x - 3| = -6}{-3} = \frac{\cancel{-3}|x - 3| = \cancel{-6}\rightarrow 2 }{\cancel{-3}} \] which leaves us with \[\large |x- 3| = 2\] Now we know that absolute values change negative numbers into positive...so in a way to show this...we break the equation we have now, into 2 equations...because if we look at that equation we have now...we can either get 2....or 02...and they will both be correct since the 2 would be changed to a 2 First equation: \[\large x - 3 = 2\] Second equation: \[\large x - 3 = -2\] And solve for 'x' in both cases

OpenStudy (anonymous):

so would be B.) ?

OpenStudy (johnweldon1993):

Correct indeed :)

OpenStudy (anonymous):

thank you so much.

OpenStudy (johnweldon1993):

Anytime!